A 33 g of sucrose solution is dissolved 350 g of teacup of 80⁰C water. Find the % m/v of

solution. Assume that the density of solution is 1.12 g/mL
1
Expert's answer
2021-09-17T02:12:56-0400

Density of solution"=\frac{Total mass}{Volume}=\frac{M}{V}"


"M=33g+350g=383g"


"\frac{383}{V}=1.12"


"V=\frac{383}{1.12}=341.96"


"\%of\frac{M}{V}=\frac{383}{341.96}×100=112\%"